OpenStudy (anonymous):

chris is 2 years old and jenny is 4years, in how many years will jen be 7 years younger than twice Chris's age?

OpenStudy (ikram002p):

chris is 2 and jenny is 4 then jenny is 2 yrs elder than chris so if chris is x age then jenny iz x+2 x+7=2(x+2)

OpenStudy (ikram002p):

x+7=2x+4 x=3 lol this one is correct

OpenStudy (ikram002p):

so henny will be 5 and chris will be 3

OpenStudy (aum):

Chris is 2, Jenny is 4. After "x" years, Chris will be (x+2) and Jenny will be (4+x) "... in how many years will jen be 7 years younger than twice Chris's age?" 4+x = 2(x+2) - 7 4+x = 2x + 4 - 7 x = 2x - 7 x = 7

OpenStudy (aum):

To check if this is correct, after 7 years, Chris will be 9, Jenny will be 11 11 = 2*9 - 7 which is true.

OpenStudy (ikram002p):

so i guess there is trivial solution ^^

OpenStudy (ikram002p):

in 1 yr also in 7 yr